OpenStudy (lilpac):

What is the probability of flipping a coin 15 times and getting heads 7 times? Round your answer to the nearest tenth of a percent. A. 19.6% B. 15.3% C. 4.2% D. 9.2%

OpenStudy (peachescherry14):

what grade is this

OpenStudy (peachescherry14):

@lilpac

OpenStudy (lilpac):

12th

OpenStudy (knov):



OpenStudy (kropot72):

This is a binomial distribution. \[\large P(7\ heads)=\frac{15!}{7!8!}\times(0.5)^{7}\times(0.5)^{8}\]